Circular No. 7001

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 K. Hurley, Space Sciences Laboratory, on behalf of the Ulysses Gamma-

Ray Burst Team; C. Kouveliotou, Universities Space Research Association (USRA); T. Murakami, Institute of Space and Astronautical Science; and T. Strohmayer, Goddard Space Flight Center (GSFC), report on behalf of a larger collaboration: “We have derived a preliminary triangulation error box for SGR 1900+14 using bursts observed in common with Ulysses and BATSE/CGRO from May through July. The corners of the box are (equinox 2000.0): α = 19^{h}07^{m}16^{s}, δ = +9^{o}23′.8; 19^{h}06^{m}52^{s}, +9^{o}06′.5; 19^{h}07^{m}45^{s}, +9^{o}39′.1; 19^{h}07^{m}21^{s}, +9^{o}21′.9. This error box includes the ROSAT source previously suggested as the quiescent x-ray counterpart to the soft γ-ray repeater (Hurley et al. 1996, Ap.J. 463, L13). We observed SGR 1900+14 with the ASCA satellite in April, and find that it has a hard spectrum consistent with a power law with index 2.2, and flux 1.3 × 10^{−11} erg cm^{−2} s^{−1} (2–10 keV). We detect coherent pulsations from the source at a high significance with a period of 5.16 s. The similarity of the period of this source to the one of 7.47 s detected from SGR 1806–20 (Kouveliotou et al. 1998, Nature 393, 235), indicates that the newly detected pulsar is the soft γ-ray repeater. The centroid of the error box of the source is at α = 19^{h}07^{m}14^{s}, δ = +9^{o}19′.3, close to the supernova remnant G42.8+0.6. As the source is still active, we urge monitoring at other wavelengths.”

C. Kouveliotou, USRA; T. Strohmayer, GSFC; K. Hurley, University of California at Berkeley; J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville (UAH); and P. Woods, UAH, report on behalf of a larger collaboration: “We initiated Target of Opportunity Observations of SGR 1900+14 with the RXTE/PCA after the source triggered BATSE/CGRO on May 30 (IAUC 6929). Our observations took place between June 2.604 and 9.045, during which we obtained 41 700 s on source. We confirm the pulsations detected with ASCA (see item above); a comparison of the ASCA and the PCA data establishes that the period is increasing at a rate of Ṗ = 6 × 10^{−11} s/s. The period and period derivative of SGR 1900+14 are very similar to those detected for SGR 1806–20 and strongly suggest that SGR 1900+14 is also a magnetar with B = 5 × 10^{14} Gauss and characteristic age P/Ṗ ∼ 1500 yr. We strongly encourage further infrared and radio observations of the source.”

1998 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green